OpenStudy (anonymous):

A worker drops a wrench from the top of a tower 62.9 m tall. What is the velocity when the wrench strikes the ground? The acceleration of gravity is 9.81 m/s^2

OpenStudy (joannablackwelder):

v^2 = vi^2 +2ax where v is final velocity, vi is initial velocity, a is acceleration, and x is displacement.

OpenStudy (joannablackwelder):

Solve for v.

OpenStudy (anonymous):

Yes. You can view it as kinetic energy gained = potential energy lost 1/2 m v^2 = m g h g = gravity and h = height from which it fell. This is equivalent to Joanna's equation, just explained a bit more.

OpenStudy (anonymous):

PE=KE, Mgh=\[Mgh=\frac{ 1 }{ 2 }Mv ^{2} =>v=\sqrt{2gh}=\sqrt{2*9.81*62.9}=42.69m/s \]

OpenStudy (anonymous):

Use Joanna's eqn, set \[v _{i} = 0\], finish it off!